OpenStudy (anonymous):

2x - 8y =5 -6x + 24y = 15

OpenStudy (anonymous):

2x = 5 + 8y -6(5 + 8y) + 24y = 15 -30 + -48y + 24y = 15 -20 + -24y = 15 -24y = 35 y = -35/24 2x = 5 + 8(-35/24) 2x = 5 + -35/3 2x = -50/3 x = -50/6 x = -25/

OpenStudy (anonymous):

okay I see what you did, thanks